OpenStudy (anonymous):

Given the circle having a diameter with endpoints (-2, 1) and (6, 7). The equation would be (x-2)2 + (y-4)2 =

OpenStudy (anonymous):

@timo86m cann you help me?

OpenStudy (anonymous):

is it 25?

OpenStudy (anonymous):

well use distance formula but use half of it tho. like this

OpenStudy (anonymous):

sqrt((6+2)^2+(7-1)^2) = 10 so half of that is 5 radious is 5 :)

OpenStudy (anonymous):

(x-2)2 + (y-4)2 =5^2

OpenStudy (anonymous):

-.- yu so smart :Dthanks!!

OpenStudy (anonymous):

(x-2)2 + (y-4)2 =5^2 is also (x-2)2 + (y-4)2 =25 :)

OpenStudy (anonymous):

aww *blush* I am not smart you are :)

OpenStudy (anonymous):

nahhh haha. omg iknew it was 25 ^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

yaha you are smart and it was 25 but it is preferable to write it (x-2)2 + (y-4)2 =5^2

OpenStudy (anonymous):

same thingggg ;D kay thank uu!